OpenStudy (anonymous):

y=ln(arcsinx)(x2-5x)1/2

OpenStudy (anonymous):

What about it?

OpenStudy (anonymous):

am asked to derive

OpenStudy (anonymous):

Okay. So is it: \[y = \ln(\sin^{-1}(x))\sqrt{x^2+5x}\]?

OpenStudy (anonymous):

ln(arcsinx)

OpenStudy (anonymous):

yeah that's the same as what I have. So we start by noting the chain rule and product rule will both rear their heads.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Therefore we need: \[\frac{d}{dx}(fg) = f'g + fg'; \frac{d}{dx}f(g(x)) = f'(g(x)) g'(x)\] So we have: \[f(x) = \ln(\arcsin(x)); g(x) = \sqrt{x^2-5x}\] Applying the product rule: \[\frac{d}{dx}f(x)g(x) = (\ln(\arcsin(x))'*\sqrt{x^2-5x} + \ln(\arcsin(x)) *(\sqrt{x^2-5x})'\] So what is: \[\frac{d}{dx}\ln(\arcsin(x)); \text{and}; \frac{d}{dx}\sqrt{x^2-5x} \text{ ?}\]

OpenStudy (anonymous):

(1/arcsinx)(1/(1-x2)......1 (2x-5).1/2(x2-5x)-1/2........2

OpenStudy (anonymous):

thank you very much

OpenStudy (anonymous):

\[ \frac{ 1 }{ arcosin(x) }\frac{ 1 }{ \sqrt{1-x^2} }...1\] \[\frac{ 2x-5 }{ 2\sqrt{x^2-5x} }...2\]

OpenStudy (anonymous):

Then do the substitutions as he said and that's it

OpenStudy (anonymous):

thank you sir